OpenStudy (beccab003):

Help with geometric mean with roots. Medal to best answer. Thanks! question will be posted below.

OpenStudy (beccab003):

If the geometric mean of a and 30 is \[4\sqrt{15}\] find the value of a

OpenStudy (anonymous):

Im read to help you! :)

OpenStudy (beccab003):

I know how to do these problems. Just not with square roots. Thanks!!!! :)

OpenStudy (anonymous):

well.. is that 4 divided by 15?

OpenStudy (anonymous):

wait nevermind

OpenStudy (anonymous):

whats your choices?

OpenStudy (anonymous):

\[\sqrt{30a}=4\sqrt{15}\]solve for \(a\)

OpenStudy (anonymous):

Dis dude always taking my questions! Im out.. Tag me if you need help on anything else

OpenStudy (beccab003):

Thanks @satellite73 but I still don't know how to continue the problem....with the roots.

Directrix (directrix):

First up, do you know the meaning of geometric mean? (If I am inadvertently butting into someone else's work, I will get out of the way.)

OpenStudy (beccab003):

the geometric mean has something to do with the nth root it think.

Directrix (directrix):

If x is the geometric mean of positive numbers a and b, then the following is true: a is to x as x is to b.

Directrix (directrix):